# Teaching Theorizing in Software Engineering Research

### v1

@@ Abstract
This chapter seeks to support software engineering (SE) researchers and educators in teaching the importance of theory as well as the theorizing process. Drawing on insights from other fields, the chapter presents 12 intermediate products of theorizing and what they mean in an SE context. These intermediate products serve different roles: some are theory products to frame research studies, some are theory generators, and others are components of theory. Whereas the SE domain doesn’t have many theories of its own, these intermediate products of theorizing can be found widely. The chapter aims to help readers to recognize these intermediate products, their role, and how they can help in the theorizing process within SE research. To illustrate their utility, the chapter then applies the set of intermediate theorizing products to the software architecture research field. The chapter ends with a suggested structure for a 12-week course on theorizing in SE which can be readily adapted by educators.
@@ 2.2 What Do Theories Look Like?
Most discussions of ‘theory’ in SE provide tutorial-style explanations suggesting that a theory consists of theoretical concepts and the relationships between them (propositions or hypotheses) (cf. stol2015theory ). This common format is a template for what is called a variance theory: if A causes B, then A and B are the concepts, and the relationship is a causal one. [^5] Most work tends to shy away from the verb ‘cause’ and use the more vaguely stated “is associated with” so as to not claim causality. Causality can typically not be demonstrated without conducting controlled experiments (see also Juristo’s chapter on experimentation in this volume), though there are some techniques that can provide convincing evidence for causality. This means that if A goes up, then B goes up also. Figure 1 presents an example of a variance theory. Variance theories are intuitive, and they can be tested using quantitative datasets, which are relatively easy to come by, for example by mining software repositories (see Codabux’s chapter in this volume) or developer surveys, notwithstanding the many challenges in both MSR and conducting surveys (see Kalinowski’s chapter in this volume).
@@
Another type of theory is called ‘process theory,’ defined by Huber and Van de Ven huber1995 as: “An explanation of the temporal order in which a discrete set of events occurred, based on a story or historical narrative.” Ralph defines a process theory as: “A process theory is a system of ideas intended to explain, understand (and sometimes predict) how an entity changes and develops,” and discusses four ‘ideal types’ of process theories frequently found in organizational/management research, and how they apply to SE ralph2018toward . Process theories are frequently more flexible and capture a wide range of phenomena in the real world. However, these theories are not easily tested using quantitative datasets only, but rather in comparison to an alternative process theory (cf. Ralph ralph2016software ). Figure 2 presents a process theory of large-scale agile software development. This process theory explains how tensions that arise through both internal and external perturbations require an emergent response of the organization or teams involved, so as to restore a certain ‘balance.’ As such, this can be seen as a dialectic process theory (one of the four ‘ideal’ types), in which “stability and change are explained by reference to the balance of power between opposing entities” van1995explaining . In this case, the opposing entities are the tensions that arise due to a conflict between agile processes and the large-scale nature of a software project, and the emergent responses by means of adopting both plan-driven and novel agile practices that can reduce those tensions. Note that this process theory defines a number of propositions (numbered P1 to P7), and may resemble the hypotheses in the variance theory in Figure 1, but that these are not easily testable using quantitative datasets.
@@ 2.3 Challenges in Developing Theory in Software Engineering Research
  • There is conflation and lack of understanding of terminology.
@@
Another major impediment to developing theory is the time commitment that is necessary. Two related factors play a role which go against this time commitment. First, there is the SE field’s pace: new technologies and trends emerge quickly. At the time of writing, the release of ChatGPT, a publicly available LLM immediately demanded attention from many researchers leading to papers that describe this new type of technology. Second, SE researchers are not rewarded for long-term investment; it is still common for researchers to target the International Conference on Software Engineering (ICSE) with papers that are not further developed. [^6] Without wishing to sound cynical or overly critical, I am not the first to lament this offutt2011purpose ; parnas2007stop ; briand2012embracing . It appears that the community norm isn’t such that researchers are encouraged or expected to follow up, nor to develop theory. This community norm is difficult to change. Students may not be overly concerned with this, as their funding clock is ticking. Supervisors are in a difficult position, because they prefer their students to succeed and not subject them to higher standards than would be the norm within the community, while both funding agencies and tenure evaluation committees are expecting results from principal investigators.
@@ 4.4 Model
Consistent with Whittle et al.’s description, the term ‘model’ usually implies that it represents an abstraction from something in the real world, and in the context of this chapter that is how we use the term model. Abstraction implies simplification and the leaving out of details. Curran curran787 , discussing statistical modeling, describes the importance of simplification:
@@
Models provide the researcher to describe something in a systematic way so as to communicate the important characteristics, leaving out the idiosyncratic issues from the real world. As an abstract representation, researchers can reason about it, perform analyses, and even computer simulations which themselves rely on simplifications stol2018abc . Simulations are discussed in more detail in De França’s chapter in this volume.
@@ 5 Components of Theory
The third set of theorizing products are actual components, or elements, of theory. These include the things we wish to reason about or measure (concepts, constructs), and the relationships between those things (statements, hypotheses). These components are commonly explained in a tutorial on ‘theory’ stol2015theory . Because the terms concept and construct are so closely related, in defining them I make reference to the other. Table LABEL:tab:components presents a summary.
@@ 5.1 Concept
Not all concepts are measurable, and therefore it is not always possible to define meaningful constructs for a concept. A concept can be any “idea that receives a name,” which also includes software development process or software product, and yet we would not seek to measure these concepts. Instead, we would measure their properties (dubin1978theory, , p. 40); for example, rather than measuring the software product, we’d seek to measure its various qualities. In other words, concepts can be much more generic ideas. Constructs are discussed in more detail below (Sec. 5.2); that section also describes one historical reason as to why the terms concept and construct may be conflated, or even considered equivalent by some.
@@ 5.2 Construct
When we adopt a quantitative research strategy, it is useful to draw the distinction, and to clarify why, I make a brief, though technical diversion. I do this by drawing on Rigdon’s explanation of what Bagozzi and Philips called the ‘Holistic Construal’ bagozzi1982representing ; rigdon2012rethinking .
@@ 5.4 Hypothesis
Sometimes the term ‘hypothesis’ is used in a more casual way, as if it were offered at a cocktail party. For example, closely related to Conway’s Law (discussed earlier) is the “Mirroring Hypothesis,” which can be stated as: “Loosely-coupled organizations will tend to develop products with more modular architectures than tightly-coupled organizations” maccormack2012exploring ; colfer2016mirroring . The form of this statement is more like a proposition than a hypothesis, because it only references concepts such as “loosely-coupled organizations” and “modular architectures.” To actually test this hypothesis, we’d need empirical indicators (constructs); how would one measure the modularity of an architecture? For example, as part of their study, MacCormack et al. operationalize a dependency between modules as a function call; if a module A calls a function in module B, then A depends on B.
@@
Statement A description relating to one or more concepts that is derived from logical reasoning, prior theoretical insights or literature, or empirical evidence. Includes propositions, typologies, and taxonomies. Statements include propositions (e.g. crowston2004effective , typologies and taxonomies ralph2018toward ; for example, Lehman’s SPE classification Lehman1980 .
@@ 6 Theorizing Products in Action: The Case of Software Architecture
The theorizing products presented above are drawn from Information Systems scholars hassan2022 . One could question, or even object, that such a theoretical focus is not applicable to a practical field like SE, a field that some perceive as an engineering field. To highlight the value of these theorizing products, I now demonstrate how theorizing takes place within the software architecture field, an area of research that is firmly rooted within the SE domain. I adopt Hassan et al.’s approach hassan2022 to illustrate how studies move the field forward by leveraging one theorizing product, to move to another theorizing product. The presentation below should not be construed as a complete and necessarily complete presentation of how theorizing actually took place. Further, while the various studies all relate to the software architecture field, the studies used to illustrate the method link to different topics, including software architecture as a concept and architectural knowledge. Reflecting on how an area of research evolved using the 12 theorizing products as a lens can help to increase awareness of theorizing so that future researchers can also learn how to reflect on a field.
Figure 6: Products of theorizing in the study of Software Architecture
@@ 6.2 From Metaphor to Analogy to Concept
Thus, an analogy was drawn to the high-level structure of a system, as suggested by the metaphors representing the architectural styles, similar to idea that one can observe different architectural styles in buildings throughout the ages, and across the world. The term ‘architecture’ is a metaphor191919‘Architecture’ is the metaphor, not the term ‘software architecture,’ because prior to the coining of the term, it didn’t exist. The same applies to ‘technical debt’: debt is the metaphor. for high-level design of a software system, and quickly found its way through into software companies, for example, in the naming of the role software architect. [^20] The analogy between building systems and construction has not gone without critique baragry2001we . I agree that the metaphor is not perfect: software architects would normally have a strong development background, whereas builders don’t usually become architects. Another point where the analogy with construction breaks down is the fact that architects in construction do, in fact, focus on details, aesthetics, and elegance. For example, architects would design ‘hidden’ gutters and downpipes as these essential elements would take away from the aesthetics of a building when visible. By now, the term ‘architecture’ (in the context of software systems) has now become so common, it is a ‘dead’ metaphor giles2001missing .
@@ 6.5 Shift in Paradigm Leads to Shift in Questions
As a result, researchers started asking different types of questions. Instead of asking, how can we capture different architectural styles or patterns in UML, for example, researchers now were focusing on defining architectural knowledge deboer2007 , and how it could be managed.
@@ 6.7 From Model to Framework
Whereas the extended model captures the full life-cycle of architecture design during which architectural knowledge (AK) can be created and used, Tang et al. tang2010comparative then turn their focus on the activities that produce and consume AK. In so doing, they identify 12 different types of activities through which producers and consumers generate and use AK.

### v2

@@ Abstract
This chapter seeks to support software engineering (SE) researchers and educators in teaching the importance of theory as well as the theorizing process. Drawing on insights from other fields, the chapter presents 12 intermediate products of theorizing and what they mean in an SE context. These intermediate products serve different roles: some are theory products to frame research studies, some are theory generators, and others are components of a theory. Whereas the SE domain doesn’t have many theories of its own, these intermediate products of theorizing can be found widely. The chapter aims to help readers to recognize these intermediate products, their role, and how they can help in the theorizing process within SE research. To illustrate their utility, the chapter then applies the set of intermediate theorizing products to the software architecture research field. The chapter ends with a suggested structure for a 12-week course on theorizing in SE which can be readily adapted by educators.
@@ 2.2 What Do Theories Look Like?
Most discussions of ‘theory’ in SE provide tutorial-style explanations suggesting that a theory consists of theoretical concepts and the relationships between them (propositions or hypotheses) (cf. stol2015theory ). This common format is a template for what is called a variance theory: if A causes B, then A and B are the concepts, and the relationship is a causal one. [^5] Most work tends to shy away from the verb ‘cause’ and use the more vaguely stated “is associated with” so as to not claim causality. Causality can typically not be demonstrated without conducting controlled experiments (see also Vegas and Juristo’s chapter on experimentation in this volume), though there are some techniques that can provide convincing evidence for causality. This means that if A goes up, then B goes up also. Figure 1 presents an example of a variance theory. Variance theories are intuitive, and they can be tested using quantitative datasets, which are relatively easy to come by, for example by mining software repositories (see Codabux et al.’s chapter in this volume) or developer surveys, notwithstanding the many challenges in both MSR and conducting surveys (see Kalinowski et al.’s chapter in this volume).
@@
Another type of theory is called ‘process theory,’ defined by Huber and Van de Ven huber1995 as: “An explanation of the temporal order in which a discrete set of events occurred, based on a story or historical narrative.” Ralph defines a process theory as: “A process theory is a system of ideas intended to explain, understand (and sometimes predict) how an entity changes and develops,” and discusses four ‘ideal types’ of process theories frequently found in organizational/management research, and how they apply to SE ralph2018toward . Process theories are frequently more flexible and capture a wide range of phenomena in the real world. However, these theories are not easily tested using quantitative datasets only, but rather in comparison to an alternative process theory (cf. Ralph ralph2016software ). Figure 2 presents a process theory of large-scale agile software development. This process theory explains how tensions that arise through both internal and external perturbations require an emergent response of the organization or teams involved, so as to restore a certain ‘balance.’ As such, this can be seen as a dialectic process theory (one of the four ‘ideal’ types), in which “stability and change are explained by reference to the balance of power between opposing entities” van1995explaining . In this case, the opposing entities are the tensions that arise due to a conflict between agile processes and the large-scale nature of a software project, and the emergent responses by means of adopting both plan-driven and novel agile practices that can reduce those tensions. Note that this process theory defines a number of propositions (numbered P1 to P7), and may resemble the hypotheses in the variance theory in Fig. 1, but that these are not easily testable using quantitative datasets.
@@ 2.3 Challenges in Developing Theory in Software Engineering Research
  • Researchers conflate and lack understanding of terminology.
@@
Another major impediment to developing theory is the time commitment that is necessary. Two related factors play a role which go against this time commitment. First, there is the SE field’s pace: new technologies and trends emerge quickly. At the time of writing, the release of ChatGPT, a publicly available LLM, immediately demanded attention from many researchers leading to papers that describe this new type of technology. Second, SE researchers are not rewarded for long-term investment; it is still common for researchers to target the International Conference on Software Engineering (ICSE) with papers that are not further developed. [^6] Without wishing to sound cynical or overly critical, I am not the first to lament this offutt2011purpose ; parnas2007stop ; briand2012embracing . It appears that the community norm isn’t such that researchers are encouraged or expected to follow up, nor to develop theory. This community norm is difficult to change. Students may not be overly concerned with this, as their funding clock is ticking. Supervisors are in a difficult position, because they prefer their students to succeed and not subject them to higher standards than would be the norm within the community, while both funding agencies and tenure evaluation committees are expecting results from principal investigators.
@@ 4.4 Model
Consistent with Whittle et al.’s description, the term ‘model’ usually implies that it represents an abstraction from something in the real world, and in the context of this chapter that is how I use the term model. Abstraction implies simplification and the leaving out of details. Curran curran787 , discussing statistical modeling, describes the importance of simplification:
@@
Models provide the researcher to describe something in a systematic way so as to communicate the important characteristics, leaving out the idiosyncratic issues from the real world. As an abstract representation, researchers can reason about it, perform analyses, and even run computer simulations which themselves rely on simplifications stol2018abc . Simulations are discussed in more detail in De França’s chapter in this volume.
@@ 5 Components of Theory
The third set of theorizing products are actual components, or elements, of theory. These include the things we wish to reason about or measure (concepts, constructs), and the relationships between those things (statements, hypotheses). These components are commonly explained in a tutorial on ‘theory’ stol2015theory . Because the terms concept and construct are so closely related, in defining one I make reference to the other. Table LABEL:tab:components presents a summary.
@@ 5.1 Concept
Not all concepts are measurable, and therefore it is not always possible to define meaningful constructs for a concept. A concept can be any “idea that receives a name,” which also includes ‘software development process’ or ‘software product,’ and yet we would not seek to measure these concepts. Instead, we would measure their properties (dubin1978theory, , p. 40); for example, rather than measuring the software product, we’d seek to measure its various qualities. In other words, concepts can be much more generic ideas. Constructs are discussed in more detail below (Sec. 5.2); that section also describes one historical reason as to why the terms concept and construct may be conflated, or even considered equivalent by some.
@@ 5.2 Construct
When we adopt a quantitative research strategy, it is useful to draw the distinction between “construct” and “concept”; to clarify why, I make a brief, though technical, diversion. I do this by drawing on Rigdon’s explanation of what Bagozzi and Philips called the ‘Holistic Construal’ bagozzi1982representing ; rigdon2012rethinking .
@@ 5.4 Hypothesis
Sometimes the term ‘hypothesis’ is used in a more casual way, as if it were offered at a cocktail party. For example, closely related to Conway’s Law (discussed earlier) is the “Mirroring Hypothesis,” which can be stated as: “Loosely-coupled organizations will tend to develop products with more modular architectures than tightly-coupled organizations” maccormack2012exploring ; colfer2016mirroring . The form of this statement is more like a proposition than a hypothesis, because it only references concepts such as “loosely-coupled organizations” and “modular architectures.” To actually test this hypothesis, we’d need empirical indicators (constructs); how would one measure the modularity of an architecture? For example, as part of their study, MacCormack et al. maccormack2012exploring operationalize a dependency between modules as a function call; if a module A calls a function in module B, then A depends on B.
@@
Statement A description relating to one or more concepts that is derived from logical reasoning, prior theoretical insights or literature, or empirical evidence. Includes propositions, typologies, and taxonomies. Statements include propositions (e.g. Crowston’s propositions on work practices in open source software crowston2004effective ), and typologies and taxonomies ralph2018toward (e.g. Lehman’s SPE classification Lehman1980 .)
@@ 6 Theorizing Products in Action: The Case of Software Architecture
The theorizing products presented above are drawn from Information Systems scholars hassan2022 . One could question, or even object, that such a theoretical focus is not applicable to a practical field like SE, a field that some perceive as an engineering field. To highlight the value of these theorizing products, I now demonstrate how theorizing takes place within the software architecture field, an area of research that is firmly rooted within the SE domain. I adopt Hassan et al.’s approach hassan2022 to illustrate how studies move the field forward by leveraging one theorizing product, to move to another theorizing product (Fig. 6). The presentation below should not be construed as a complete and necessarily accurate presentation of how theorizing actually took place. Further, while the various studies all relate to the software architecture field, the studies used to illustrate the method link to different topics, including software architecture as a concept and architectural knowledge. Reflecting on how an area of research evolved using the 12 theorizing products as a lens can help to increase awareness of theorizing so that future researchers can also learn how to reflect on a field.
Figure 6: Example map of products of theorizing in the study of software architecture. This map is only one illustrative example of the author’s (possibly naive) interpretation of how studies relate to one another. The author acknowledges this map is incomplete. Further, in creating a map, any theorizing product can be linked to any other. Creating a more complete map can be a useful yet time-consuming way to learn about a research area
@@ 6.2 From Metaphor to Analogy to Concept
Thus, an analogy was drawn to the high-level structure of a system, as suggested by the metaphors representing the architectural styles, similar to the idea that one can observe different architectural styles in buildings throughout the ages, and across the world. The term ‘architecture’ is a metaphor191919‘Architecture’ is the metaphor, not the term ‘software architecture,’ because prior to the coining of the term, it didn’t exist. The same applies to ‘technical debt’: debt is the metaphor. for high-level design of a software system, and quickly found its way through into software companies, for example, in the naming of the role software architect. [^20] The analogy between building systems and construction has not gone without critique baragry2001we . I agree that the metaphor is not perfect: software architects would normally have a strong development background, whereas builders don’t usually become architects. Another point where the analogy with construction breaks down is the fact that architects in construction do, in fact, focus on details, aesthetics, and elegance. For example, architects would design ‘hidden’ gutters and downpipes as these essential elements would take away from the aesthetics of a building when visible. By now, the term ‘architecture’ (in the context of software systems) has become so common, it is a ‘dead’ metaphor giles2001missing .
@@ 6.5 Shift in Paradigm Leads to Shift in Questions
As a result, researchers started asking different types of questions. Instead of asking how we can capture different architectural styles or patterns in UML, for example, researchers now were focusing on defining architectural knowledge deboer2007 , and how it could be managed.
@@ 6.7 From Model to Framework
Whereas the extended model captures the full life-cycle of architecture design during which architectural knowledge (AK) can be created and used, Tang et al. tang2010comparative then turned their focus on the activities that produce and consume AK. In so doing, they identify 12 different types of activities through which producers and consumers generate and use AK.